# Efficient Graph Optimization via Distance-Aware Graph Representation Learning

### v6

@@ A.6 Stability Analysis
Stability Proof.
The stability follows from the Lipschitz continuity of the DRTR representation function and the bounded influence of edge changes on the overall representation.
@@
Let $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ be the edge sets of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, respectively, with $|\mathcal{E}_{1}\triangle\mathcal{E}_{2}|\leq\Delta$.
The change in representation can be bounded by:
$\|\mathbf{Z}_{1}-\mathbf{Z}_{2}\|_{F}\leq\sum_{k=1}^{K}\gamma_{k}\|\mathbf{H}_{1}^{(k)}-\mathbf{H}_{2}^{(k)}\|_{F}.$
For each hop $k$, the change in aggregation is bounded by the number of affected edges:
$\|\mathbf{H}_{1}^{(k)}-\mathbf{H}_{2}^{(k)}\|_{F}\leq L_{k}\cdot\Delta\cdot\sqrt{|\mathcal{V}|},$
where $L_{k}$ is the Lipschitz constant for hop $k$.
Combining these bounds gives the desired stability result. ∎
A.7 Additional Optimization Properties
A.7.1 Convexity Analysis
Proposition 2 (Convexity of DRTR Components).
The following components of the DRTR objective are convex:
  • $\mathcal{L}_{\text{regularization}}(\theta)$ is strongly convex
  • $\mathcal{L}_{\text{DR}}(\theta)$ is convex in the distance parameters
@@ A.7.1 Convexity Analysis
  • The overall objective is non-convex due to the classification loss
A.7.2 Saddle Point Analysis
Theorem 7 (Saddle Point Avoidance).
@@ A.7.2 Saddle Point Analysis
Under mild conditions on the initialization and learning rate, the DRTR algorithm avoids saddle points with high probability.
Saddle Point Proof.
The proof relies on the fact that the DRTR objective has a non-zero gradient almost everywhere, and the stochastic nature of the optimization helps escape saddle points.
The key insight is that the distance recomputation and topology reconstruction components introduce sufficient randomness to avoid getting stuck in saddle points. ∎
A.8 Hyperparameter Sensitivity Analysis
A.8.1 Learning Rate Schedule
The optimal learning rate schedule for DRTR follows:
$\eta_{t}=\frac{\eta_{0}}{\sqrt{t}}\cdot\exp(-\mu t),$
@@ A.8.1 Learning Rate Schedule
where $\eta_{0}$ is the initial learning rate and $\mu$ controls the decay rate.
A.8.2 Regularization Parameters
The regularization parameters $\lambda_{1},\lambda_{2},\lambda_{3}$ should satisfy:
$\lambda_{1}+\lambda_{2}+\lambda_{3}=1,\quad\lambda_{i}\geq 0\text{ for }i=1,2,3.$
The optimal values can be found through cross-validation or adaptive methods.
A.9 Extension to Other Graph Learning Tasks
A.9.1 Link Prediction
For link prediction tasks, the DRTR framework can be extended by:
  • Using the learned representations to compute edge probabilities
  • Incorporating structural features in the similarity computation
  • Adapting the topology reconstruction for link-specific objectives
A.9.2 Graph Classification
For graph-level tasks, DRTR can be extended by:
  • Aggregating node representations using graph-level pooling
  • Incorporating graph-level structural features
  • Adapting the distance recomputation for graph-level objectives
A.10 Implementation Details
A.10.1 Numerical Stability
To ensure numerical stability, we implement:
  • Gradient clipping to prevent exploding gradients
  • Layer normalization to stabilize training
  • Temperature scaling for attention computation
A.10.2 Efficient Computation
For large-scale graphs, we use:
  • Sparse matrix operations for adjacency matrices
  • Batch processing for distance computation
  • Approximate nearest neighbor search for topology reconstruction
A.11 Comprehensive Visualization Analysis
In this section, we provide comprehensive visualizations that illustrate the theoretical properties, optimization behavior, and empirical performance of the DRTR framework.
A.11.1 Performance Comparison: Radar Chart Analysis
Figure 3: Radar chart comparing DRTR variants against baseline models across multiple performance metrics. The chart demonstrates DRTR’s superior performance in accuracy, stability, and robustness while maintaining competitive efficiency and scalability.
Figure 3 presents a comprehensive radar chart analysis comparing different GNN models across five key metrics: Accuracy, Stability, Efficiency, Robustness, and Scalability. The visualization clearly demonstrates that DRTR-enhanced models (particularly GKHDDRA) achieve superior performance across most metrics, with significant improvements in stability and robustness.
A.11.2 Optimization Landscape: 3D Heatmap Visualization
Figure 4: 3D visualization of the DRTR optimization landscape showing the loss function surface and convergence path. The plot illustrates the smooth optimization trajectory and absence of sharp local minima.
Figure 4 provides a 3D visualization of the DRTR optimization landscape, illustrating the smooth nature of the loss function and the convergence path. The visualization demonstrates that DRTR avoids sharp local minima and exhibits stable convergence behavior, supporting our theoretical convergence guarantees.
A.11.3 Convergence Analysis: Spindle Plot
Figure 5: Spindle plot showing convergence behavior and performance distribution over time. Left: Convergence curves with confidence intervals. Right: Box plots showing performance distribution across different time windows.
Figure 5 presents a spindle plot analysis of convergence behavior, combining convergence curves with performance distribution analysis. The left panel shows the convergence trajectories with confidence intervals, while the right panel displays box plots illustrating performance distribution across different time windows. This visualization confirms DRTR’s faster and more stable convergence compared to baseline methods.
A.11.4 Embedding Visualization: t-SNE Analysis
Figure 6: t-SNE visualization of node embeddings learned by DRTR. Left: Ground truth cluster structure. Right: DRTR embeddings showing improved cluster separation and reduced noise.
Figure 6 demonstrates the quality of node embeddings learned by DRTR through t-SNE visualization. The comparison between ground truth clusters and DRTR embeddings shows that our method successfully learns meaningful representations with improved cluster separation and reduced noise, validating the effectiveness of the distance recomputation and topology reconstruction mechanisms.
A.11.5 Complexity Analysis: Scalability Visualization
Figure 7: Comprehensive complexity analysis showing time complexity, space complexity, scalability, and efficiency trade-offs. The plots demonstrate DRTR’s favorable complexity characteristics with minimal overhead.
Figure 7 provides a comprehensive complexity analysis across four dimensions: time complexity, space complexity, scalability, and efficiency trade-offs. The visualizations confirm that DRTR maintains competitive computational complexity while achieving superior performance, supporting our theoretical complexity bounds.
A.11.6 Theoretical Validation: Convergence and Generalization
Figure 8: Theoretical analysis plots validating convergence rates and generalization bounds. Left: Convergence rate comparison showing DRTR’s faster convergence. Right: Generalization bound comparison demonstrating tighter bounds for DRTR.
Figure 8 validates our theoretical results through empirical visualization. The left panel demonstrates DRTR’s superior convergence rate compared to standard optimization methods, while the right panel shows tighter generalization bounds, confirming the theoretical advantages of our approach.

### v7

@@ A.7 Stability Analysis
Proof A.13 (Stability Proof).
@@
By Lipschitz continuity and bounded edge influence, with $|\mathcal{E}_{1}\triangle\mathcal{E}_{2}|\leq\Delta$:
$\|\mathbf{Z}_{1}-\mathbf{Z}_{2}\|_{F}\leq\sum_{k=1}^{K}\gamma_{k}\|\mathbf{H}_{1}^{(k)}-\mathbf{H}_{2}^{(k)}\|_{F}\leq\sum_{k=1}^{K}\gamma_{k}L_{k}\cdot\Delta\cdot\sqrt{|\mathcal{V}|}=C\cdot\Delta\cdot\sqrt{|\mathcal{V}|},$
where $C=\sum_{k=1}^{K}\gamma_{k}L_{k}$ and $L_{k}$ is the Lipschitz constant for hop $k$.
A.8 Additional Optimization Properties
Proposition A.14 (Convexity).
@@ A.8 Additional Optimization Properties
$\mathcal{L}_{\text{regularization}}$ is strongly convex, $\mathcal{L}_{\text{DR}}$ is convex in distance parameters, while the overall objective is non-convex due to classification loss.
Proposition A.15 (Saddle Point Avoidance).
@@
Under mild conditions, DRTR avoids saddle points with high probability due to non-zero gradients and stochastic optimization.
A.9 Hyperparameter Optimization
The learning rate schedule:
$\eta_{t}=\frac{\eta_{0}}{\sqrt{t}}\cdot\exp(-\mu t),$ (59)
@@ A.9 Hyperparameter Optimization
where $\eta_{0}$ is initial rate and $\mu$ controls decay. Regularization parameters satisfy:
$\lambda_{1}+\lambda_{2}+\lambda_{3}=1,\quad\lambda_{i}\geq 0.$ (60)
A.10 Extension to Other Graph Learning Tasks
DRTR extends to link prediction and graph classification by adapting similarity computation and pooling operations. For link prediction, edge probabilities are computed from learned representations using the topology reconstruction similarity function. For graph classification, node representations are aggregated via graph-level pooling (mean/attention/hierarchical), with distance recomputation adapted for inter-graph similarities.
A.11 Implementation Details
A.11.1 Layer Normalization Details
Layer normalization is applied after each hop aggregation to ensure numerical stability and prevent vanishing gradients. The detailed formulation is:
where $\mu_{v}^{(k)}=\frac{1}{d}\sum_{i=1}^{d}h_{v,i}^{(k)}$ is the mean, $\sigma_{v}^{(k)}=\sqrt{\frac{1}{d}\sum_{i=1}^{d}(h_{v,i}^{(k)}-\mu_{v}^{(k)})^{2}}$ is the standard deviation, and $\epsilon=10^{-5}$ is a small constant to prevent division by zero.
A.11.2 Numerical Stability
We implement gradient clipping ($\|\nabla\mathcal{L}\|_{2}\leq\tau_{\text{clip}}$), layer normalization (Eq. (4)), and temperature scaling ($\tau_{k}=\tau_{0}\exp(-\eta k)$) to ensure numerical stability during training.
A.11.3 Computational Optimization
Sparse matrix operations reduce memory from $\mathcal{O}(|\mathcal{V}|^{2})$ to $\mathcal{O}(|\mathcal{E}|)$. Batch processing optimizes distance computation:
$T_{\text{batch}}=\left\lceil\frac{|\mathcal{V}|}{B}\right\rceil\cdot\mathcal{O}(B\cdot k\cdot d)=\mathcal{O}(|\mathcal{V}|\cdot k\cdot d),$ (62)
where $B$ is batch size. Approximate $k$-NN search reduces TR complexity from $\mathcal{O}(|\mathcal{V}|^{2}\cdot d)$ to $\mathcal{O}(|\mathcal{V}|\cdot k\cdot d)$.
A.11.4 Implementation Settings
Table LABEL:tab:implementation-settings summarizes the key hyperparameters and implementation settings used in our DRTR framework across different datasets.
[…]
tab:implementation-settings Setting Category Parameter Value Training Learning Rate 0.005 Weight Decay 0.0005 Epochs 1000 Early Stopping Patience 100 Batch Size Full graph Numerical Stability Gradient Clipping Threshold 1.0 Layer Normalization $\epsilon$ $10^{-5}$ Initial Temperature $\tau_{0}$ 1.0 Temperature Decay Rate $\eta$ 0.1 Distance Recomputator Initial Scaling $\lambda_{0}$ 0.1 Scaling Decay Rate $\rho$ 0.05 Minimum Scaling $\lambda_{\min}$ 0.01 Pruning Percentile $p$ 0.75 Topology Reconstructor Similarity Threshold $\beta$ 0.5 Edge Addition Probability $\theta$ 0.6 Approximate NN Search (k) 50 Multi-Hop Aggregation Maximum Hops $K$ 3 Hidden Dimension $d^{\prime}$ 64 Attention Head Dimension 8 Computation Sparse Matrix Format CSR Distance Batch Size 1024
Table 5: Implementation settings and hyperparameters for DRTR framework.